OpenStudy (anonymous):

A polyhedron has 10 faces and 24 edges. What is the number of vertices of the polyhedron?

Parth (parthkohli):

Euclid's Formula states that,\[E + 2 = F + V\]

Parth (parthkohli):

It's easy to plug-in \(E = 24\) and \(F = 10\). We have the following equation to solve:\[24 + 2 = 10 + V\]

Parth (parthkohli):

@Shiv23 can you do the rest? Just asking, class 8 mein ho kya?

Parth (parthkohli):

Callisto returns!

OpenStudy (anonymous):

is it 16

Parth (parthkohli):

That's right.

OpenStudy (anonymous):

thnx

Parth (parthkohli):

My pleasure :)